OpenStudy (anonymous):

hat is the sum of a 22–term arithmetic sequence where the first term is 55 and the last term is –29

OpenStudy (anonymous):

22/2(55+(-29)) = 286

OpenStudy (anonymous):

Sn=22/2(55-29)=286